Solve the differential equation by Bernoulli equation. Show complete solution.


dy/dx + 1/3 y = 1/3 (1 + 3x) y4

1
Expert's answer
2021-10-06T18:09:05-0400
"y'+\dfrac{1}{3}y=\dfrac{1}{3}(1+3x)y^4"

"v=y^{1-n}=y^{1-4}=y^{-3}"

"v'=-3y^{-4}y'"

"-\dfrac{1}{3}v'+\dfrac{1}{3}v=\dfrac{1}{3}(1+3x)"

"v'-v=-1-3x"

Integrating factor


"\mu(x)=e^{-x}"

"e^{-x}v'-e^{-x}v=-e^{-x}(1+3x)"

"d(e^{-x}v)=-e^{-x}(1+3x)dx"

Integrate


"\int d(e^{-x}v)=-e^{-x}(1+3x)dx"

"\int xe^{-x}dx"

"\int udv=uv-\int vdu"

"u=x, du=dx"

"dv=e^{-x}dx, v=\int e^{-x}dx=-e^{-x}"

"\int xe^{-x}dx=-xe^{-x}+\int e^{-x}dx"

"=-xe^{-x}-e^{-x}+C_1"


"e^{-x}v=e^{-x}+3xe^{-x}+3e^{-x}+C"

"v=4+3x+Ce^x"

"y=\dfrac{1}{\sqrt[3]{4+3x+Ce^x}}"